# The Effect of the Gotthard Base Tunnel on Road Traffic: A Synthetic Control Approach

### v1

@@ 1 Introduction
The rationale for the construction of an efficient railway infrastructure differs across countries. For Switzerland, investments in a flat, energy-efficient, and high-performance railway infrastructure are conceived as an important contribution aimed at limiting the negative environmental impact on the sensitive alpine landscape of road transport, especially freight transportation (Albalate and Bel, 2012). However, robust ex-post analyses of the impact of new rail infrastructure on induced demand and modal shift from private car to rail remain scarce. To the best of our knowledge, only the study of Borsati and Albalate (2020) systematically examines the causal impact of high-performance railway infrastructure on motorway, concluding that neither the opening of a high-speed rail (HSR) nor the opening of on-track competition led to a modal shift from motorway to HSR services. Moreover, the review of Givoni and Dobruszkes (2013) on HSR indicates that the demand for HSR several years after inauguration is about ten to twenty percent induced demand. The rest is attributed to mode substitution, whereas, in most cases, most HSR passengers have traveled with conventional rail before. In contrast, the substitution from aircraft, car, and coach is usually modest. However, there exist varying effects (Givoni and Dobruszkes, 2013).
@@
These findings underscore the need for a new, rigorous and empirically driven evaluation of infrastructure policies. To fill this research gap, we focus on the Gotthard Base Tunnel, a part of the New Rail Link through the Alps. The construction of the Gotthard Base Tunnel took 17 years and was completed in December 2016, with a total cost of 12.2 billion Swiss francs. In our study, we answer the research question of whether the travel time savings due new Gotthard Base Tunnel in the short run reduced the number of light vehicles on the parallel motorway section. Therefore, knowing the average car occupancy rate in leisure travel, we can also approximate whether the new connection induced an increase in passengers.
@@
The case of the Gotthard Base Tunnel is interesting for several reasons. While most research has focused on inter-modal competition between HSR and air services, especially on long point-to-point links between cities, our case of the Gotthard Base Tunnel differs in relevant aspects. The section with the Gotthard Base Tunnel from Art-Goldau to Bellinzona is relatively short, with a travel time of 58 minutes, which means that the private car is a relevant competitor. Such studies are relatively scarce (see for an example the study of Borsati and Albalate, 2020). Most researched HSR connections have a length of several hundred kilometers (Cascetta, Papola, Pagliara, and Marzano, 2011, Cheng, 2010, Givoni and Dobruszkes, 2013). Second, Switzerland has no HSR trains, only an intercity network. Moreover, the Swiss rail system is “open”, which means that customers can board a train at any time without a reservation. Fourth, there are no special prices for the Gotthard Base Tunnel connection. Together with the fact that Switzerland has a dense suburban public transport service to every municipality, this eliminates the problem of many of the HSR examples studied being difficult for customers to access or having no connection to the countryside at the destination (Sánchez-Mateos and Givoni, 2012). Another distinguishing feature is the touristic use of the route, including “transit traffic”; for example, tourists from Germany travel to Italy. A recent study has shown that 97 percent of these people travel by private car and only 3 percent by train (Ohnmacht, Hüsser, Balthasar, Liebrich, Geoffroy, Reckermann, Bolenz, Kowald, et al., 2024).
@@
In our comparative case study, we apply the synthetic control (Abadie, Diamond, and Hainmueller, 2010) and the synthetic difference in differences (Arkhangelsky, Athey, Hirshberg, Imbens, and Wager, 2021) methods to construct a synthetic Gotthard motorway section—a counterfactual that estimates the number of vehicles that would have used the Gotthard alpine crossing in the absence of the tunnel. The synthetic Gotthard motorway section is a weighted average of other alpine crossings in Switzerland and neighboring countries, serving as the donor pool. Our results indicate that, on average, the number of light vehicles on the Gotthard motorway section decreased by 164 or 152 per day, depending on the method, during the months of April to October. This—using the numbers of leisure travelers—corresponds to around 300 fewer people traveling by car through the Gotthard every day. These effects correspond to a 1% reduction compared to the pre-treatment average observed in the years prior to the opening of the Gotthard Base Tunnel. However, despite the modest size of the effect, the 95% bootstrap confidence intervals amount to [-265; -89] and [-291; -84], with all values below zero. Sensitivity analyses when controlling for potential spill over effects confirm the robustness of this small but statistically significant effect on the parallel motorway section. Moreover, when comparing the Gotthard motorway section with other Swiss motorway sections prone to leisure traffic, the negative effect on daily vehicles increases to 610. However, this effect is not statistically significant. When comparing these numbers with the simultaneous daily increase of approximately 2,000 rail passengers (e.g., from 2016 to 2017), we conclude that induced traffic is the main part of the additional passengers on the Gotthard rail route.
@@
Our paper proceeds as follows. In Section 2, we discuss the related literature for our case study. Section 3 introduces the Gotthard Base Tunnel in greater detail. In Section 4, we present our study design and the method used to estimate the causal effect of interest. Section 5 shows the data and first descriptive results. In Section 6, we present the main results and those of the sensitivity analyses. Section 7 discusses the results and concludes.
@@ 2 Literature review
In the context of tourism, for travel distances up to 400 km, the main distinction in travel mode choice is between the car and public transport, while for journeys exceeding 400 km, air transport becomes the dominant mode (Thrane, 2015). Regarding mode choice between car and public transport—as with mode choice between other transport modes—one can differentiate between push measures that discourage car use and pull measures that encourage public transport use (see, e.g., Zarabi, Waygood, Olsson, Friman, and Gousse-Lessard, 2024). However, a research gap exists concerning hard pull measures that intervene at the level of physical infrastructure. It is well known that demand, measured for example by the number of guests, increases when infrastructure is improved, as shown by, e.g., Pagliara, La Pietra, Gomez, and Vassallo (2015). Also, several studies show that demand for public transport increases due to large-scale rail investments, e.g., the high-speed rail (HSR) in Europe increased inter-city traffic between major metropolitan areas in North-west Europe (see, e.g., Vickerman, 2015). According to Givoni and Dobruszkes (2013), between 75% and 90% of new HSR demand is drawn from other transport modes. However, the extent and source of modal shift vary significantly across routes, depending on factors such as route distance, number of stops, travel time, and the speed of the HSR service. Depending on these characteristics, the shift may primarily occur from air travel, conventional rail, coach, or private car. HSR services tend to achieve the highest market shares when in-vehicle travel time ranges between 1 and 3.5 hours (Givoni and Dobruszkes, 2013).
@@ 3 Background
The Swiss political system operates as a direct democracy. On September 27, 1992, Swiss voters approved the New Rail Link through the Alps [^4] In German: Neue Eisenbahn-Alpentransversale (NEAT). (NRLA) initiative, with 64 percent voting in favor. The NRLA represents Switzerland’s largest construction project in its history. [^5] See https://www.eda.admin.ch/aboutswitzerland/en/home/swiss-stories/bauwerk-im-dienste-europas.html (accessed on December 4, 2024). It includes three tunnels: the Lötschberg Tunnel (34.6 km), the Gotthard Base Tunnel (57.1 km)—the longest railway tunnel in the world—, and the Ceneri Base Tunnel (15.4 km). Switzerland invested approximately 22.8 billion Swiss Francs in the construction of the NRLA, nearly 3.5 % of its Gross Domestic Product (GDP). The NRLA facilitates the operation of longer and heavier trains, allowing more freight trains to traverse the Alps with fewer locomotives and in less time. This enhancement not only accelerates passenger travel times but also offers a wider range of connections within Switzerland and across Europe. Moreover, the Gotthard route’s rail transport capacity has greatly increased: up to 260 freight trains and 65 passenger trains can now travel daily at speeds up to 250 km/h (whereas in reality, due to efficiency reasons, they usually travel daily at speeds up to 200 km/h). With the opening of the Gotthard Base Tunnel, there was initially no increase in the number of trains. An (major) expansion of rail services only occurred following the commissioning of the Ceneri Base Tunnel.
@@
The modal split in passenger transport along the Gotthard axis after the commissioning of the Gotthard Base Tunnel shifted towards public transport usage (for a descriptive analysis see ARE, 2023). The demand for public transportation increased by about 4,000 passengers per day from 2013 to 2019, rising from about 8,500 to 12,000, an increase of 41%. From 2016 to 2017, the daily number of passengers increased from about 9,100 to 11,200, an increase of around 22%. Figure 2 shows that after 2016—i.e., following the opening of the Gotthard Base Tunnel—the number of rail passengers on the Gotthard route increased more rapidly than the average number of passengers on long-distance routes of the Swiss Federal Railways. By 2019, passenger numbers on the Gotthard route had reached 141% of their 2013 level, whereas the national average had increased by only about 16% over the same period. Moreover, in Figure 2, we also see that only about 5% of the passengers used 2016 the Gotthard mountain route, whereas 95% traveled through the Gotthard Base Tunnel.
@@ 4 Identification and estimation
We aim to assess the effect of the Gotthard Base Tunnel on the parallel motorway section, i.e., number of vehicles in the Gotthard road tunnel, which represents a single treated unit. While we can observe the vehicles in the Gotthard road tunnel, identifying the impact of the Gotthard Base Tunnel requires a comparison with a counterfactual—a scenario of the Gotthard motorway section to show what would have happened if the Gotthard Base Tunnel had not occurred. To construct this scenario, we use the synthetic control method (see, e.g., Abadie, Diamond, and Hainmueller, 2010, Abadie, 2021), selecting control units from other alpine crossings that did not receive comparable major rail transport infrastructure investments. These alpine crossings form a so-called donor pool, allowing us to create a ’synthetic Gotthard motorway section’ that represents what might have occurred in on the Gotthard motorway section without the Gotthard Base Tunnel. This synthetic version is derived by assigning weights to each alpine crossing in the donor pool based on their similarity to the Gotthard alpine crossing prior to the Gotthard Base Tunnel’s opening, aiming to minimize pre-treatment differences.
In this study, we employ a panel data set consisting of $n$ alpine crossings, indexed by $i\in\{1,...,n\}$, observed over $T$ time periods, each denoted by $t\in\{1,...,T\}$. The variable $Y_{it}$ represents the observed outcome (i.e., number of vehicles) for unit $i$ in period $t$. The unit $i=n$ (the Gotthard motorway section) experiences the effect of the new rail transport infrastructure starting in the post-treatment period $T_{0}+1$. According to Huber (2023), we calculate the estimated effect of the intervention, $\hat{\Delta}_{n,T=t}$, for the Gotthard in the periods following the intervention ($t\geq T_{0}+1$). This estimate is computed as the difference between the observed outcome on the Gotthard motorway section and a weighted average of outcomes from the non-intervened units—other alpine crossings—during the same periods. The synthetic control estimator $\hat{\Delta}_{n,T=t}$ is thus defined as:
@@
The sum of all weights for the non-treated units sum up to one, $\sum_{i=1}^{n-1}\hat{w}_{i}=1$, and each weight $\hat{w}_{i}$ is non-negative ($\hat{w}_{i}\geq 0$).
In addition to the synthetic control method, we also employ more recent synthetic difference-in-differences approach as suggested by Arkhangelsky, Athey, Hirshberg, Imbens, and Wager (2021).
Ensuring that our estimates identify the machanism of interest, i.e., the effect of the Gotthard Base tunnel on the vehicles on the parallel motorway section relies on assumptions about the data-generating process (see, e.g., Huntington-Klein (2021)). In the following, we present the identifying assumptions underlying our analysis (see also, e.g., Abadie (2021) or more recently in the transport literature Wallimann (2024)):
@@
Assumption 4 implies that there must be no spillover effects of the Gotthard Base tunnel on other alpine crossings. We discuss Assumption 4 in greater detail in Section 6.2.
@@
To implement the synthetic control method using the statistical software R, we utilize the functions available in the synth package as detailed by Hainmueller and Diamond (2015). Additionally, we apply the synthetic difference-in-differences method through the synthdid package, provided by Arkhagelsky (2021).

### v2

@@
Acknowledgments: We are indebted to Antonin Danalet for helpful comments. Moreover, Zeliya Schär provided excellent research assistance.
@@ 1 Introduction
The rationale for the construction of an efficient railway infrastructure differs across countries. For Switzerland, investments in a flat, energy-efficient, and high-performance railway infrastructure are conceived as an important contribution aimed at limiting the negative environmental impact on the sensitive alpine landscape of road transport (Albalate and Bel, 2012). However, robust ex-post analyses of the impact of new rail infrastructure on induced demand and modal shift from private car to rail remain scarce. To the best of our knowledge, only the study of Borsati and Albalate (2020) systematically examines the causal impact of high-performance railway infrastructure on motorway, concluding that neither the opening of a high-speed rail (HSR) nor the opening of on-track competition led to a modal shift from motorway to HSR services. Moreover, the review of Givoni and Dobruszkes (2013) on HSR indicates that the demand for HSR several years after inauguration is about ten to twenty percent induced demand. The rest is attributed to mode substitution, whereas, in most cases, most HSR passengers have traveled with conventional rail before. In contrast, the substitution from aircraft, car, and coach is usually modest, although the effects vary across contexts and studies (Givoni and Dobruszkes, 2013).
@@
These findings underscore the need for a new, rigorous and empirically driven evaluation of infrastructure policies. To fill this research gap, we focus on the Gotthard Base Tunnel, a part of the New Rail Link through the Alps. The construction of the Gotthard Base Tunnel took 17 years and was completed in December 2016, at a total cost of 12.2 billion Swiss francs. In our study, we answer the research question of whether the travel time savings due to the new Gotthard Base Tunnel in the short run reduced the number of light vehicles on the parallel motorway section. Therefore, knowing the average car occupancy rate in leisure travel, we can also approximate whether the new connection induced an increase in passengers.
@@
The case of the Gotthard Base Tunnel is interesting for several reasons. While most research has focused on inter-modal competition between HSR and air services, especially on long point-to-point links between cities, our case of the Gotthard Base Tunnel differs in relevant aspects. The section with the Gotthard Base Tunnel from Art-Goldau to Bellinzona is relatively short, with a travel time of 58 minutes, which means that the private car is a relevant competitor. Such studies are relatively scarce (see for an example the study of Borsati and Albalate, 2020). Most researched HSR connections have a length of several hundred kilometers (Cascetta, Papola, Pagliara, and Marzano, 2011, Cheng, 2010, Givoni and Dobruszkes, 2013). Second, Switzerland has no HSR trains, only an intercity network. Third, the Swiss rail system is “open”, which means that customers can board a train at any time without a reservation. Fourth, there are no special prices for the Gotthard Base Tunnel connection. Together with the fact that Switzerland has a dense suburban public transport service to every municipality, this eliminates the problem of many of the HSR examples studied being difficult for customers to access or having no connection to the countryside at the destination (Sánchez-Mateos and Givoni, 2012). Another distinguishing feature is the touristic use of the route, including “transit traffic”; for example, tourists from Germany travel to Italy through the Gotthard axis, crossing the Swiss Alps. A recent study has shown that 97 percent of these people travel by private car and only 3 percent by train (Ohnmacht, Hüsser, Balthasar, Liebrich, Geoffroy, Reckermann, Bolenz, Kowald, et al., 2024).
@@
In our comparative case study, we apply the synthetic control (Abadie, Diamond, and Hainmueller, 2010) and the synthetic difference in differences (Arkhangelsky, Athey, Hirshberg, Imbens, and Wager, 2021) methods to construct a synthetic Gotthard motorway section—a counterfactual that estimates the number of vehicles that would have used the Gotthard alpine crossing in the absence of the tunnel. The synthetic Gotthard motorway section is a weighted average of other alpine crossings in Switzerland and neighboring countries, serving as the donor pool. Our results indicate that, on average, the number of light vehicles on the Gotthard motorway section decreased by 135 or 152 per day, depending on the method, during the months of April to October. This—using the numbers of leisure travelers—corresponds to around 300 fewer people traveling by car through the Gotthard every day. These effects correspond to a just under 1% reduction compared to the pre-treatment average observed in the years prior to the opening of the Gotthard Base Tunnel. However, despite the modest size of the effect, the 95% bootstrap confidence intervals amount to [-269; -127] and [-291; -84], with all values below zero. Sensitivity analyses, e.g., when controlling for potential spillover effects, confirm the robustness of this small but statistically significant effect on the parallel motorway section. Moreover, when comparing the Gotthard motorway section with other Swiss motorway sections prone to leisure traffic—instead of other alpine crossings—, the negative effect on daily vehicles increases to 617. However, this effect is not statistically significant. When comparing these numbers with the simultaneous daily increase of approximately 2,000 rail passengers (e.g., from 2016 to 2017), we conclude that induced traffic is the main part of the additional passengers on the Gotthard rail route.
@@
Our paper proceeds as follows. In Section 2, we discuss the related literature for our case study. Section 3 introduces the Gotthard Base Tunnel in greater detail. In Section 4, we present our research design, outlining both the identification strategy and the econometric method used to estimate the causal effect of interest. Section 5 shows the data and first descriptive results. In Section 6, we present the main results and those of the sensitivity analyses. Section 7 discusses the results and concludes.
@@ 2 Literature review
In the context of tourism, for travel distances up to 400 km, the main distinction in travel mode choice is between the car and public transport, while for journeys exceeding 400 km, air transport becomes the dominant mode (Thrane, 2015). Regarding mode choice between car and public transport—as with mode choice between other transport modes—one can differentiate between push measures that discourage car use and pull measures that encourage public transport use (see, e.g., Zarabi, Waygood, Olsson, Friman, and Gousse-Lessard, 2024). However, a research gap exists concerning hard pull measures that intervene at the level of physical infrastructure. It is well known that demand, measured, for example, by the number of travelers, increases when infrastructure is improved, as shown by, e.g., Pagliara, La Pietra, Gomez, and Vassallo (2015). Also, several studies show that demand for public transport increases due to large-scale rail investments, e.g., the high-speed rail (HSR) in Europe increased inter-city traffic between major metropolitan areas in North-west Europe (see, e.g., Vickerman, 2015). According to Givoni and Dobruszkes (2013), between 75% and 90% of new HSR demand is drawn from other transport modes. However, the extent and source of modal shift vary significantly across routes, depending on factors such as route distance, number of stops, travel time, and the speed of the HSR service. Depending on these characteristics, the shift may primarily occur from air travel, conventional rail, coach, or private car. HSR services tend to achieve the highest market shares when in-vehicle travel time ranges between 1 and 3.5 hours (Givoni and Dobruszkes, 2013).
@@ 3 Background
The Swiss political system operates as a direct democracy. On September 27, 1992, Swiss voters approved the New Rail Link through the Alps [^4] In German: Neue Eisenbahn-Alpentransversale (NEAT). (NRLA) initiative, with 64 percent voting in favor. The NRLA represents Switzerland’s largest construction project in its history. [^5] See https://www.eda.admin.ch/aboutswitzerland/en/home/swiss-stories/bauwerk-im-dienste-europas.html (accessed on December 4, 2024). It includes three tunnels: the Lötschberg Base Tunnel (34.6 km), the Gotthard Base Tunnel (57.1 km)—the longest railway tunnel in the world—, and the Ceneri Base Tunnel (15.4 km). Switzerland invested approximately 22.8 billion Swiss Francs in the construction of the NRLA, nearly 3.5 % of its Gross Domestic Product (GDP). The Lötschberg Base Tunnel went into operation in 2007.666The number of public transport passengers increased by 59% within one year (see, e.g., Danalet, 2025).
The NRLA facilitates the operation of longer and heavier trains, allowing more freight trains to traverse the Alps with fewer locomotives and in less time. This enhancement not only accelerates passenger travel times but also offers a wider range of connections within Switzerland and across Europe. Moreover, the Gotthard route’s rail transport capacity has greatly increased: up to 260 freight trains and 65 passenger trains can now travel daily at speeds up to 250 km/h (whereas in reality, due to efficiency reasons, they usually travel daily at speeds up to 200 km/h). With the opening of the Gotthard Base Tunnel, there was initially no increase in the number of trains. An (major) expansion of rail services only occurred following the commissioning of the Ceneri Base Tunnel.
@@
The modal split in passenger transport along the Gotthard axis after the commissioning of the Gotthard Base Tunnel shifted towards public transport usage (for a descriptive analysis see ARE, 2023). The demand for public transportation increased by about 3,500 passengers per day from 2013 to 2019, rising from about 8,500 to 12,000, an increase of 41%. From 2016 to 2017, the daily number of passengers increased from about 9,100 to 11,200, an increase of around 22%. Figure 2 shows that after 2016—i.e., following the opening of the Gotthard Base Tunnel—the number of rail passengers on the Gotthard route increased more rapidly than the average number of passengers on long-distance routes of the Swiss Federal Railways. By 2019, passenger numbers on the Gotthard route had reached 141% of their 2013 level, whereas the national average had increased by only about 16% over the same period. Moreover, in Figure 2, we also see that only about 5% of the passengers used 2016 the Gotthard mountain route, whereas 95% traveled through the Gotthard Base Tunnel.
@@ 4 Identification and estimation
4.1 Methodology and implementation
We aim to assess the effect of the Gotthard Base Tunnel on the parallel motorway section, i.e., the number of vehicles in the Gotthard road tunnel, which represents a single treated unit. While we can observe the vehicles in the Gotthard road tunnel, identifying the impact of the Gotthard Base Tunnel requires a comparison with a counterfactual—a scenario of the Gotthard motorway section to show what would have happened if the Gotthard Base Tunnel had not occurred. To construct this scenario, we use the synthetic control method (see, e.g., Abadie, Diamond, and Hainmueller, 2010, Abadie, 2021), selecting control units from other alpine crossings that did not receive comparable major rail transport infrastructure investments. These alpine crossings form a so-called donor pool, allowing us to create a ’synthetic Gotthard motorway section’ that represents what might have occurred in on the Gotthard motorway section without the Gotthard Base Tunnel. This synthetic version is derived by assigning weights to each alpine crossing in the donor pool based on their similarity to the Gotthard alpine crossing prior to the Gotthard Base Tunnel’s opening, aiming to minimize pre-treatment differences.
In this study, we employ a panel data set consisting of $n$ alpine crossings, indexed by $i\in\{1,...,n\}$, observed over $T$ time periods, each denoted by $t\in\{1,...,T\}$. The variable $Y_{it}$ represents the observed outcome (i.e., number of vehicles) for unit $i$ in period $t$. The unit $i=n$ (the Gotthard motorway section) experiences the effect of the new rail transport infrastructure starting in the post-treatment period $T_{0}+1$, with $T_{0}$ denoting the last pre-treatment period, i.e., 2016. According to Huber (2023), we calculate the estimated effect of the intervention, $\hat{\Delta}_{n,T=t}$, for the Gotthard in the periods following the intervention ($t\geq T_{0}+1$). This estimate is computed as the difference between the observed outcome on the Gotthard motorway section and a weighted average of outcomes from the non-intervened units—other alpine crossings—during the same periods. The synthetic control estimator $\hat{\Delta}_{n,T=t}$ is thus defined as:
@@ 4.1 Methodology and implementation
The synthetic control method selects a weight vector $W=(\hat{w}_{1},\dots,\hat{w}_{n-1})$, with $\hat{w}_{i}\geq 0$ and $\sum_{i=1}^{n-1}\hat{w}_{i}=1$. That means that the sum of all weights for the non-treated units sum up to one and are non-negative. The goal is to minimize the squared difference between the treated unit’s outcome and the weighted combination of control units over the pre-treatment period.
In addition to the synthetic control method, we also employ more recent synthetic difference-in-differences (SDID) approach as suggested by Arkhangelsky, Athey, Hirshberg, Imbens, and Wager (2021).
To implement the synthetic control method using the statistical software R, we utilize the functions available in the synth package as detailed by Hainmueller and Diamond (2015), allowing to apply the synthetic control group method for comparative case studies. as described by Abadie and Gardeazabal (2003) and Abadie, Diamond, and Hainmueller (2010). Additionally, we apply the synthetic difference-in-differences method through the synthdid package, provided by Arkhagelsky (2021), which allows the implementation of the SDID estimator for the average treatment effect in panel data, as proposed by Arkhangelsky, Athey, Hirshberg, Imbens, and Wager (2019).
4.2 Assumptions
Ensuring that our estimates identify the mechanism of interest, i.e., the effect of the Gotthard Base tunnel on the vehicles on the parallel motorway section, relies on assumptions about the data-generating process (see, e.g., Huntington-Klein (2021)). In the following, we present the identifying assumptions underlying our analysis (see also, e.g., Abadie (2021) or more recently in the transport literature Wallimann (2024)):
@@ 4.2 Assumptions
Assumption 4 implies that there must be no spillover effects of the Gotthard Base tunnel on other alpine crossings. However, it is conceivable that, for example, a traveler journeying from Hamburg to Milan might alter her route if less congestion is expected on the Gotthard route. For instance, on April 30, 2025, at 11 p.m., travel via the Gotthard takes according to Google Maps 11 hours and 57 minutes, while using the San Bernardino Tunnel takes 11 hours and 48 inutes—making the two routes nearly identical in terms of travel time. Nevertheless, during peak hours, the Gotthard motorway section may now be somewhat less congested, meaning that travelers, after the tunnel’s introduction, may opt not to switch to the San Bernardino Tunnel. Similar spillover effects might occur for other journeys. Therefore, we assess the sensitivity of our results regarding Assumption 4 in greater detail in Section 6.2.